A student buys a TV set of R11000 on a hire purchase agreement. He agrees to pay a deposit of 20% and the balance over a period of two years in a monthly installment at 14% simple interest p. a. He also takes out and insurance premium of R18,50 per month. Determine the amount of each installment


1
Expert's answer
2021-03-10T07:07:33-0500

Hire purchase =(deposit +Monthly installment)

Deposit =(20/100)*11000

=R2200

Remaining Amount = 11000-2200=8800

A=P(1+in)

Subitute the values

Amount=8800(1+(0.14*2))

 =R11264

Months =2*12

=24Months

Therefore =11264/24

=R469.33

Monthly insurance =1850

Total monthly insurance =1850+469

=R2319